### v2

# Quadrature Over-the-Air-Computing for Multimodal Dual-Stream Signal Processing

Abstract
We propose a novel quadrature over-the-air computing (Q-OTAC) framework that enables the simultaneously computation of two independent functions and/or data streams within a single transmission. In contrast to conventional OTAC schemes, where a single function is computed by treating each complex signal as a single component, the proposed Q-OTAC exploits both in-phase and quadrature (IQ) components of a complex signal, encoding two distinct functions and/or data streams at the edge devices (EDs) and employing a novel low-complexity IQ-decoupled combiner at the access point (AP) to independently recover each stream, which effectively doubles the computation rate. A key strength of this framework lies in its simplicity and broad compatibility: the extension into the quadrature domain is conceptually straightforward, yet remakably powerful, allowing seamless integration into existing OTAC techniques. Simulation results validate the effectiveness of this approach, including the first demonstration of dual-function aggregation (e.g., parallel summation and product), highlighting the potential of Q-OTAC for enabling multi-modal and high-efficiency beyond fifth generation (B5G) applications.
Index Terms—  Quadrature, over-the-air computing, OTAC, AirComp, multifunctionality, multimodal.
1 Introduction
The expected convergence of communication and distributed computing functionalities in B5G/6G wireless networks [1, 2], has fueled the development of function-centric communication paradigms, including the increasingly prominent over-the-air computing (OTAC) [3, 4, 5], also referred to as AirComp. Rather than transmitting individual data streams for centralized processing, OTAC allows multiple edge devices to simultaneously transmit pre-processed signals that are aggregated over the wireless medium – leveraging the inherent superposition property of the wireless multiple-access channel – effectively computing a global function directly at the physical layer, via a wide class of nomographic functions, including summations, averages, and geometric means [6, 7].
Driven by the growing demand for low-latency, energy-efficient, and bandwidth-constrained applications, such as federated learning and sensor fusion in autonomous networks and massive internet of things (IoT), OTAC is emerging as a key enabler of integrated communication and computing (ICC) in future wireless systems [8, 9]. Furthermore, its relevance becomes more prominent in a broader context when considering the integration of computation with integrated sensing and communications (ISAC) [10, 11] – towards integrated sensing, communication, and computing (ISCC)– where concurrent communications, sensing, and signal processing over wireless channels will be critical to achieve real-time intelligence and system efficiency [12, 13, 14].
However, despite significant research progress - including broadband channel adaptations [15], digital combining [16], and correlation handling [17] - one of the challenges of OTAC is essentially related to its computational capacity. In particular, current methods allow only a single function to be computed per resource instance (e.g., one antenna, one frequency, one symbol). This limitation arises from the standard practice of mapping data onto the in-phase component of complex wireless signals, while leaving the quadrature component unused for data encoding and only for precoding. As a result, the system underutilizes the available complex signal space, and the computation rate does not fully harness its potential.
This architectural bottleneck echoes similar limitations that were overcome in other areas of wireless system design - by leveraging the inherent parallel domains of the complex field. For instance, quadrature spatial modulation (QSM) [18] extends conventional spatial modulation (SM) [19] by jointly and independently modulating the IQ components of transmitted symbols, doubling the spectral efficiency and improving performance without increasing antenna count [20].
Inspired by such developments, we propose in this article quadrature over-the-air computing (Q-OTAC), a novel framework that enables the simultaneous computation of two independent nomographic functions and/or two data symbols by leveraging the IQ domains of a single complex-valued transmission. This is enabled by encoding two streams of pre-processed symbols at the transmitter to each of the IQ domains, and applying a novel IQ-decoupled minimum mean square error (MMSE)-based combiner that independently estimates each target function at the AP.
The proposed scheme, to the best of the authors’ knowledge, is the first OTAC/AirComp method to realize dual-function aggregation per single transmission resource, achieving double the total computation rate. Moreover, the novel capacity of the Q-OTAC to support heterogeneous data inputs and/or functions makes it well-suited for multimodal aggregation tasks, relevant for applications in B5G, such as intelligent vehicular networks, edge computing, and semantic inference [21, 22].
2 System Model and Conventional OTAC
Consider a single-input multiple-output (SIMO) wireless uplink system with $K$ single-antenna EDs and a single AP equipped with $N$ receive antennas. Under perfect synchronization, the received signal vector $\mathbf{y}\in\mathbb{C}^{N\times 1}$ at the $N$ antennas of the AP is given by
$\mathbf{y}=\sum_{k=1}^{K}\mathbf{h}_{k}p_{k}s_{k}+\mathbf{w}=\mathbf{H}\mathbf{P}\mathbf{s}+\mathbf{w}\in\mathbb{C}^{N\times 1},\vskip-2.15277pt$ (1)
where $\mathbf{H}=[\mathbf{h}_{1},\ldots,\mathbf{h}_{K}]\in\mathbb{C}^{N\times K}$ is the channel matrix concatenating the $K$ ED-to-AP channel vectors $\mathbf{h}_{k}\in\mathbb{C}^{N\times 1}$, $\mathbf{P}=\mathrm{diag}(\mathbf{p})\in\mathbb{C}^{K\times K}$ is the diagonal matrix containing the local precoding weights $p_{k}\in\mathbb{C}$ of the EDs, collected as a vector in $\mathbf{p}\triangleq[p_{1},\ldots,p_{K}]^{\mathsf{T}}\in\mathbb{C}^{K\times 1}$, $\mathbf{s}=[s_{1},\ldots,s_{K}]^{\mathsf{T}}\in\mathbb{C}^{K\times 1}$ is the vector containing the transmit symbols $s_{k}$ of the $k$-th ED, and $\mathbf{w}\sim\mathcal{CN}(\mathbf{0},\sigma^{2}_{w}\mathbf{I})\in\mathbb{C}^{N\times 1}$ is the additive white Gaussian noise (AWGN) with variance $\sigma^{2}_{w}\in\mathbb{R}^{+}$.
In this work, we consider the practical assumption where feedback is unavailable between the AP and the EDs, such that the transmit precoders $p_{k}$ to be fixed and known a priori. This eliminates coordination overhead and improves robustness in dense or latency-sensitive scenarios. Therefore, in the remainder of this article, we will consider $p_{k}=1\penalty 10000\ \forall k$.
2.1 State-of-the-Art OTAC Scheme
In the conventional single-stream OTAC setting, the AP aims to compute a multivariate target function $f(d_{1},\ldots,d_{K}):\mathbb{R}^{K}\to\mathbb{R}$ of the data symbols $d_{1},\ldots,d_{K}\in\mathbb{R}$ at the $K$ EDs. The target function is assumed to be nomographic22footnotemark: 2 [6], which means that it can be decomposed as a post-processed sum of pre-processed symbols, $i.e.$,
$f(d_{1},\ldots,d_{K})\triangleq\psi\bigg(\sum_{k=1}^{K}\underbrace{\varphi_{k}(d_{k})}_{\triangleq\;s_{k}}\bigg)=\psi\bigg(\sum_{k=1}^{K}s_{k}\bigg)\in\mathbb{R},$
where $\varphi_{k}(\cdot)$ denotes the local pre-processing function at the $k$-th ED, $\psi(\cdot)$ denotes the global post-processing function at the AP, and $s_{k}\triangleq\varphi_{k}(d_{k})\in\mathbb{R}$ has been defined as the effective transmit symbol of the $k$-th ED after pre-processing - and is ultimately equivalent to the transmit symbols of eq. (1).
For example, when the target function is a summation, the pre-/post-processing functions are simply identity functions, $i.e.$, $\varphi_{k}^{\mathrm{SUM}}(x)=x$ and $\psi^{\mathrm{SUM}}(x)=x$, such that $f^{\mathrm{SUM}}(d_{1},\ldots,d_{K})\!\triangleq\!\psi^{\mathrm{SUM}}\big(\sum_{k=1}^{K}\varphi_{k}^{\mathrm{SUM}}(d_{k})\big)\!=\!\sum_{k=1}^{K}d_{k}.$
11footnotetext: Many real-valued multivariable functions used in practice – such as summation, average, weighted average, product, geometric mean, $p$-norms, and polynomials – are directly nomographic and thus well-suited for OTAC.
On the other hand, when the target function is a product, the pre-/post-processing functions are the logarithmic and exponentiation to an arbitrary base $b$, $i.e.$, $\varphi_{k}^{\mathrm{PROD}}(x)=\log_{b}(x)$ and $\psi^{\mathrm{PROD}}(x)=b^{\;\!x}=\mathrm{exp}_{b}(x)$, respectively, such that $f^{\mathrm{PROD}}(d_{1},\ldots,d_{K})\triangleq\exp_{b}\!\big(\sum_{k=1}^{K}\log_{b}(d_{k})\big)\!=\!\prod_{k=1}^{K}d_{k}.$
Then, given that the transmit symbols at the EDs are generated via pre-processing the data symbols as $s_{k}\triangleq\varphi_{k}(d_{k})\in\mathbb{R}$, the received signal model of eq. (1) can be rewritten as
$\mathbf{y}=\sum_{k=1}^{K}\mathbf{h}_{k}s_{k}+\mathbf{w}=\sum_{k=1}^{K}\mathbf{h}_{k}\varphi_{k}(d_{k})+\mathbf{w}\in\mathbb{C}^{N\times 1},\vskip-4.30554pt$ (3)
where it can be seen that the wireless multiple-access channel inherently achieves the summation of the pre-processed symbols via the superposition phenomena.
Therefore, the objective of the AP is to remove the effect of the channel to yield the summation of only the pre-processed symbols, such that the post-processing function can be applied to find the target function output. Of course, each symbol can be estimated independently then recombined to yield the desired sum, but it should be stated clearly that this is not the objective of OTAC, and instead, the objective is to apply a single combiner vector $\mathbf{u}\in\mathbb{C}^{N\times 1}$ to the received signal to obtain the estimate of the target function, $i.e.$,
$\tilde{f}(\mathbf{u};\mathbf{y})=\psi\big(\mathbf{u}^{\mathsf{T}}\mathbf{y}\big)\approx f(d_{1},\ldots,d_{K})\in\mathbb{R}.$ (4)
In all, given the channel matrix $\mathbf{H}$, the AP aims to construct an optimal combiner which minimizes the mean square error (MSE) of the estimated target function [^2] Assuming a feedback channel between the AP and the EDs, the precoding weights $p_{k}$ can also be jointly optimized with the combiner $\mathbf{u}$ in order to minimize the objective function. However, this consideration is out of scope of this article, and will be considered in an extended work., formulated as
$\mathbf{u}_{\mathrm{opt}}$ $=\operatorname*{arg\,min}_{\mathbf{u}}\,\mathbb{E}\!\left[\left|f(d_{1},\ldots,d_{K})-\psi(\mathbf{u}^{\mathsf{T}}\mathbf{y})\right|^{2}\right].\vskip-8.61108pt$ (5)
While various state-of-the-art (SotA) methods exist for designing $\mathbf{u}$ – such as branch-and-bound, iterative, and message-passing schemes – we focus on MMSE-based linear combiners due to their simplicity in admitting a closed-form solution with a strong baseline performance [23, 24], with
$\mathbf{u}_{\mathrm{MMSE}}=\big(\mathbf{H}\mathbf{H}^{\mathsf{H}}+\sigma^{2}_{w}\mathbf{I}_{N\times N}\big)^{\!-1}\!\cdot\mathbf{H}\cdot\mathbf{1}_{K\times 1}\in\mathbb{C}^{N\times 1}.\vskip-4.30554pt$ (6)
Since the combiner in eq. (6) is complex-valued, the combined output $\mathbf{u}^{\mathsf{T}}\mathbf{y}$ is also complex. However, since the original data symbols and the nomographic target functions in OTAC are real-valued, as defined in eq. (LABEL:eq:OTACtargetfunction), only the real part of the combined signal is relevant. Thus, in practice, the function estimate is obtained using the post-processing function only unto the real component of the combined output, $i.e.$,
$\tilde{f}(\mathbf{u};\mathbf{y})=\psi\!\left(\mathrm{Re}\{\mathbf{u}^{\mathsf{T}}\mathbf{y}\}\right)\in\mathbb{R}.\vskip 3.22916pt$ (7)
3 Proposed Dual-Stream Q-OTAC
As can be seen in Section 2, the conventional OTAC paradigm processes only a single real-valued function of scalar data symbols per transmission, fundamentally underutiliing the complex signal space by discarding the quadrature component, and also inherently limiting the computational capacity of each resource instance to one computing stream.
To address this limitation, we propose a dual-stream quadrature over-the-air computing (Q-OTAC) framework that encodes two real-valued data symbols – one in each of the IQ components of the complex transmit symbol – with a novel combiner design which enables an independent computation of the two nomographic functions at the AP, which has been illustrated in Fig. 1, clearly highlighting the two streams.
3.1 Proposed Q-OTAC Dual-stream Transmitter Design
First, let us denote the two independent OTAC target functions by $f(\,\cdots):\mathbb{R}^{K}\to\mathbb{R}$ and $g(\,\cdots):\mathbb{R}^{K}\to\mathbb{R}$, with the corresponding pre-/post-processing functions given by $\psi^{\mathrm{f}}(\cdot),\psi^{\mathrm{g}}(\cdot)$ and $\varphi^{\mathrm{f}}_{k}(\cdot),\varphi^{\mathrm{g}}_{k}(\cdot)$, respectively, such that
$f(d^{\mathrm{\;\!f}}_{1},\ldots,d^{\mathrm{\;\!f}}_{K})$ $\triangleq\psi^{\mathrm{f}}\Big(\textstyle\sum_{k=1}^{K}\varphi^{\mathrm{f}}_{k}(d^{\mathrm{\;\!f}}_{k})\Big)\in\mathbb{R},$ (8a)
$g(d^{\mathrm{\;\!g}}_{1},\ldots,d^{\mathrm{\;\!g}}_{K})$ $\triangleq\psi^{\mathrm{g}}\Big(\textstyle\sum_{k=1}^{K}\varphi^{\mathrm{g}}_{k}(d^{\mathrm{\;\!g}}_{k})\Big)\in\mathbb{R},\vskip-2.15277pt$ (8b)
where $d^{\mathrm{\;\!f}}_{k}\in\mathbb{R}$ and $d^{\mathrm{\;\!g}}_{k}\in\mathbb{R}$ are the two independent data stream at the $k$-th ED, respectively for each target function.
In light of the above, the Q-OTAC scheme proposes to transmit the two streams of the pre-processed data symbols $s^{\mathrm{f}}_{k}\!\triangleq\!\varphi_{k}^{\mathrm{f}}(d_{k}^{\mathrm{\;\!f}})\!\in\!\mathbb{R}$ and $s^{\mathrm{g}}_{k}\!\triangleq\!\varphi_{k}^{\mathrm{g}}(d_{k}^{\mathrm{\;\!g}})\!\in\!\mathbb{R}$ respectively in the IQ domain of the complex transmit symbol, $i.e.$,
$s_{k}\triangleq s^{\mathrm{f}}_{k}+js^{\mathrm{g}}_{k}=\varphi_{k}^{\mathrm{f}}(d_{k}^{\mathrm{\;\!f}})+j\varphi_{k}^{\mathrm{g}}(d_{k}^{\mathrm{\;\!g}})\in\mathbb{C},\vskip-2.15277pt$ (9)
where $j\triangleq\sqrt{-1}$ is the imaginary unit.
The resulting received signal, with $p_{k}=1$, is given by
$\mathbf{y}=\sum_{k=1}^{K}\mathbf{h}_{k}\big(s_{k}^{\mathrm{f}}\!+\!js_{k}^{\mathrm{g}}\big)+\mathbf{w}=\mathbf{H}(\mathbf{s}^{\mathrm{f}}\!+\!j\mathbf{s}^{\mathrm{g}})+\mathbf{w}\!\in\!\mathbb{C}^{N\times 1}\vskip-6.45831pt$ (10)
where $\mathbf{s}^{\mathrm{f}}\!\triangleq\![s^{\mathrm{f}}_{1},\ldots,s^{\mathrm{f}}_{K}]^{\mathsf{T}}\!\in\!\mathbb{R}^{N\times 1}$, $\mathbf{s}^{\mathrm{g}}\!\triangleq\![s^{\mathrm{g}}_{1},\ldots,s^{\mathrm{g}}_{K}]^{\mathsf{T}}\!\in\!\mathbb{R}^{N\times 1}$ are the vectors containing the in-phase and quadrature symbols transmitted by the EDs, respectively.
Equation (10) illustrates the key novelty of the proposed Q-OTAC framework in the complex-valued transmit symbol containing two data streams in each of the IQ components, contrasting with the conventional OTAC in eq. (3) with only a real-valued transmit symbol composed of only one stream.
Fig. 1: Illustration of the proposed Q-OTAC system.
3.2 Proposed Q-OTAC Dual-stream Combiner Design
First, the received signal model of eq. (10) is decoupled and reformulated to a real-valued equivalent system as
$\bm{y}\triangleq\begin{bmatrix}\Re\{\mathbf{y}\}\\ \Im\{\mathbf{y}\}\end{bmatrix}\!=\!\underbrace{\begin{bmatrix}\Re\{\mathbf{H}\}&\hskip-6.45831pt-\Im\{\mathbf{H}\}\\ \Im\{\mathbf{H}\}&\Re\{\mathbf{H}\}\end{bmatrix}}_{\triangleq\bm{H}\in\mathbb{R}^{2N\times 2N}}\hskip-9.68747pt\overbrace{\begin{bmatrix}\mathbf{s}^{\mathrm{f}}\\ \mathbf{s}^{\mathrm{g}}\\ \end{bmatrix}}^{\triangleq\bm{s}\in\mathbb{R}^{2N\times 1}}\hskip-9.68747pt+\!\!\underbrace{\begin{bmatrix}\Re\{\mathbf{w}\}\\ \Im\{\mathbf{w}\}\end{bmatrix}}_{\triangleq\bm{w}\in\mathbb{R}^{2N\times 1}}\!\!\in\mathbb{R}^{2N\times 1},$
where $\bm{y}\in\mathbb{R}^{2N\times 1},\bm{H}\in\mathbb{R}^{2N\times 2N},\bm{s}\in\mathbb{R}^{2N\times 1},\bm{w}\in\mathbb{R}^{2N\times 1}$ are the IQ-decoupled received signal vector, transmit signal vector and AWGN vector, respectively.
In light of the above, we aim to design two distinct combiners $\bm{u}^{\mathrm{f}}\in\mathbb{R}^{2N\times 1}$ and $\bm{u}^{\mathrm{g}}\in\mathbb{R}^{2N\times 1}$ to be applied directly unto the received signal vector $\bm{y}$}, which respectively yields the estimated target function values $\tilde{f}$ and $\tilde{g}$, $i.e.$,
$\tilde{{f}}(\bm{u}^{\mathrm{f}};\bm{y})$ $=\psi^{\mathrm{f}}\big((\bm{u}^{\mathrm{f}}){}^{\mathsf{\!T}}\,\!\bm{y}\big)\,\approx f(d^{\mathrm{\;\!f}}_{1},\ldots,d^{\mathrm{\;\!f}}_{K})\in\mathbb{R},$ (12a)
$\tilde{g}(\bm{u}^{\mathrm{g}};\bm{y})$ $=\psi^{\mathrm{g}}\big((\bm{u}^{\mathrm{g}}){}^{\mathsf{\!T}}\,\!\bm{y}\big)\approx g(d^{\mathrm{\;\!g}}_{1},\ldots,d^{\mathrm{\;\!g}}_{K})\in\mathbb{R}.$ (12b)
Given the above, the MMSE  [^3] As with conventional OTAC, other optimization methods for combiner design may be considered, however, in this work, we focus on the closed-form, low-complexity linear MMSE-based combiner as a proof of concept of the proposed dual-stream framework. problem is formulated as
$\mathbf{u}^{\mathrm{f}}_{\mathrm{opt}}$ $=\operatorname*{arg\,min}_{\mathbf{u}^{\mathrm{f}}}\,\mathbb{E}\Big[\big|f(d^{\mathrm{\;\!f}}_{1},\ldots,d^{\mathrm{\;\!f}}_{K})-\tilde{f}(\bm{u}^{\mathrm{f}};\bm{y})\big|^{2}\Big]$ (13a)
$\equiv\operatorname*{arg\,min}_{\mathbf{u}^{\mathrm{f}}}\,\mathbb{E}\Big[\big|\underbrace{\mathbf{1}_{N\times 1}^{\mathsf{T}}\!\cdot\!\mathbf{s}^{\mathrm{f}}}_{\sum_{k=1}^{K}\!s^{\mathrm{f}}_{k}}\,-\,(\bm{u}^{\mathrm{f}})^{\mathsf{\!T}}\,\!\!\cdot\!(\bm{H}\bm{s}+\bm{w})\big|^{2}\Big],$
$\mathbf{u}^{\mathrm{g}}_{\mathrm{opt}}$ $=\operatorname*{arg\,min}_{\mathbf{u}^{\mathrm{g}}}\,\mathbb{E}\Big[\big|g(d^{\mathrm{\;\!g}}_{1},\ldots,d^{\mathrm{\;\!g}}_{K})-\tilde{g}(\bm{u}^{\mathrm{g}};\bm{y})\big|^{2}\Big]$ (13b)
$\equiv\operatorname*{arg\,min}_{\mathbf{u}^{\mathrm{g}}}\,\mathbb{E}\Big[\big|\underbrace{\mathbf{1}_{N\times 1}^{\mathsf{T}}\!\cdot\!\mathbf{s}^{\mathrm{g}}}_{\sum_{k=1}^{K}\!s^{\mathrm{g}}_{k}}\,-\,(\bm{u}^{\mathrm{g}})^{\mathsf{\!T}}\,\!\!\cdot\!(\bm{H}\bm{s}+\bm{w})\big|^{2}\Big],$
where the respective inverses of the post-processing functions has been applied to the two terms of the squared error, while retaining the equivalence of the minimization problems as the post-processing function of nomographic functions are typically strictly monotonic ($i.e.$, identity, scalar division, exponentiation) such that the minimization of the squared error in the transformed domain, $i.e.$, $|\psi(b)-\psi(a)|^{2}$, is equivalent to the minimization in the original domain, $|b-a|^{2}$.
In light of the above, the closed-form solutions of the MMSE combiners of eq. (13) are efficiently obtained as
$\bm{u}_{\mathrm{MMSE}}^{\mathrm{f}}$ $=\!\big(({\bm{H}}{\bm{H}}^{\mathsf{T}}\!+\!\tfrac{\sigma^{2}_{w}}{2}\mathbf{I}_{2N\times 2N})\big)^{\!-1}{\bm{H}}\bm{c}^{\mathrm{f}}\in\mathbb{R}^{2N\times 1}\!,\!\!$ (14a)
$\bm{u}_{\mathrm{MMSE}}^{\mathrm{g}}$ $=\!\big(({\bm{H}}{\bm{H}}^{\mathsf{T}}\!+\!\tfrac{\sigma^{2}_{w}}{2}\mathbf{I}_{2N\times 2N})\big)^{\!-1}{\bm{H}}\bm{c}^{\mathrm{g}}\!\in\mathbb{R}^{2N\times 1}\!,\!\!$ (14b)
where the auxiliary domain selector vectors $\bm{c}^{\mathrm{f}}\in\mathbb{R}^{2N\times 1}$ and $\bm{c}^{\mathrm{g}}\in\mathbb{R}^{2N\times 1}$ are simply given by
$\bm{c}^{\mathrm{f}}\triangleq\begin{bmatrix}\mathbf{1}_{N\times 1}\\ \mathbf{0}_{N\times 1}\end{bmatrix}\in\mathbb{R}^{2N\times 1},\penalty 10000\ \text{and}\penalty 10000\ \bm{c}^{\mathrm{g}}\triangleq\begin{bmatrix}\mathbf{0}_{N\times 1}\\ \mathbf{1}_{N\times 1}\end{bmatrix}\in\mathbb{R}^{2N\times 1}.$ (15)
In all, by configuring the pre/post-processing functions of two independent target functions, the proposed Q-OTAC framework enables the parallel computation of two distinct target functions and/or of two distinct sets of data ($i.e.$, two independent streams), by applying two highly efficient MMSE-based combiners $\bm{c}^{\mathrm{f}}$ and $\bm{c}^{\mathrm{g}}$ to the received signal.
4 Performance Analysis
Fig. 2: The NMSE CDF for different multimodal dual-stream Q-OTAC scenarios for $N=20,K=20$ and $\mathrm{SNR}=15\mathrm{dB}$.
In this section, we demonstrate the core innovation of the proposed dual-stream Q-OTAC scheme and provide the first proof-of-concept that the OTAC framework can achieve simultaneous computation of two distinct target functions using only a single wireless transmission resource and a low-complexity (closed-form) linear combiner at the AP. Specifically, the performance of the proposed dual-stream Q-OTAC scheme is validated and analyzed via numerical simulation, comparing against the conventional single-stream OTAC baseline using the linear MMSE combiner of eq. (6).
First, Figure 2 presents the cumulative distribution functions (CDFs) of the normalized mean square error (NMSE) for various heterogeneous function pairs, including simultaneous arithmetic mean and weighted average (with arbitrary weights), as well as a simultaneous summation and product. In all cases, the proposed Q-OTAC scheme maintains reliable estimation accuracy across functions.
A relative performance variation among the different function combinations can also be observed, but this is not a remnant of the combiner design, but reflects the underlying numerical behavior of the pre-/post-processing of the nomographic functions. For example, the product function involves logarithmic and exponential operations, which are inherently more sensitive to noise than the identity mappings used for the summation and mean.
(a) CDF of the NMSE ($\mathrm{SNR=15\text{dB}}$).
(b) Average NMSE against SNR.
Fig. 3: Summation performance of the proposed Q-OTAC against the conventional OTAC, with $N=20$, $K=20$.
Next, Figure 3 compares the performance of the proposed dual-stream Q-OTAC with the conventional single-stream OTAC scheme, evaluated in terms of the total NMSE per computation via the CDF and average NMSE over SNR, in the respective subfigures. The results demonstrate that Q-OTAC achieves a noticable improved total NMSE per computation, over the entire range of both the CDF and SNR plots with an approximate gain of $\penalty 10000\ 5\text{dB}$ thanks to its novel and efficient utilization of both the IQ components of the complex signal to perform computing, as opposed to only a single real component in the SotA, thereby doubling the signal space and effective computation rate.
5 Conclusion and Future Works
We proposed the quadrature over-the-air computing (Q-OTAC), the first OTAC framework to enable dual-function aggregation in a single transmission via a novel IQ-decoupled MMSE combiner. Simulation results confirm its effectiveness in heterogeneous function settings, achieving twice the computation rate and improved average accuracy over the single-stream OTAC, supporting multimodal and scalable B5G OTAC scenarios. Future works will investigate extended combiner designs, jointly with the precoder optimization considerations, and incoporate ICC/ISCC approaches.